Startup Night 14 20140912 UT

02:17 MODS1 + LBC-R

Turn on LBC-R (only) --- I realized that you still have to toggle between pages to see the update.

In IIFGUI, pseudo-mono checked with MODS1 dominant.

On left side: acqMODS PDS456clumps.acq

On right side: dofpia, /redonly, /first:
  • not in keyhole
  • Z7 = +2500, Z8 = -500
  • didnt work
  • clear ao again
  • Z8 = +2500 --- at least coma is out.

Meanwhile, aligning MODS. On first offset, PDS looks to be well centered. Blind offset and slit image (0004)
execMODS PDS456_600sec.obs

dofpia is converging now, albeit there is a lot of spherical.

dofpia, /redonly while locating OB.

02:56 Playing PDS456i_long.ob which will use 120-sec exposures at i-SLOAN

None of the images were in focus and the tec & lbc sci images look like they are pupil images. The LBC seemed to have disconnected ? I ran the software down and back up again. Seems ok -

03:05 running dofpia again.

03:08 playing ob again (10 60-sec exposures)
  • this is not working.
    • 03158 has several jumps.
    • see guidelbc.dx.20140912lbto.tmp
    • and thumbnail image starts with star in corner and star eventually goes out of box. There are plenty of stars in the tech chip images.
03:17 unchecked blue channel/mirror (even though that was blue channel was turned off in OB, blue side was not turned on and left side was not connected).
  • trying OB again.
    • 032105 --- still shows jumps

03:21 playing ob again, this time unchecking guiding in the UI
  • now the star is in the middle of the thumbnail image (on exp 032723)
  • and the image is good.

03:22 MODS Red IMCS failed to lock. Airmass 1.752. LDGROT = 266 deg. "r"etry did the job.
  • 034323 - ok (lbc not guiding)
  • 034549 - not ok (lbc guiding)

03:51 obtained a series of images between which dithered by same amount as object.
  • not well collimated --- we're at low airmass, but I would have thought that the collimation on the science camera would be better. guide star is very close to MODS field of view.

03:54 untick guiding to see whether any effect on MODS images...
  • again, star goes to middle of thumbnail. no noticeable effect on mods images - still out of collimation.

04:06 Series of offsets in imaging mode with MODS. Thru-slit image on qso. 2 x 180-sec spec on qso.

04:25 MOS Align check: acqMODS mods.1.modsplck.acq (seeing is about 1")

Computed Mask Alignment Offset:
   dX:  -3.766 arcsec
   dY:  -1.166 arcsec
  dPA: -0.1658 degrees

MODS Offset Command:
   offsetpointing -0.1658 -3.766 -1.166 detxy rel

04:40 MOS Align check: execMODS StoneO _1136_act0_5dicerelxy.img (5dice is a misnomer). 859 ms on guider (below lower limit). seeing is 0.75".

arget:
  OBJNAME StoneO_1136
  OBJCOORDS 21:48:03.9798 +00:00:13.540
  GUINAME GStar
  GUICOORDS 21:48:05.2364 -00:03:01.170
  POSANGLE 0.0
  AGWFILT Clear
  PRESET ACTIVE

Instrument:
  InstConfig red imaging
  SlitMask Imaging

Exec:
  Object StoneO_1136
  ExpTime 30
  NImgs 1
  red filter r_sdss
  WFSWait
  red GO
  # Remaining imaging commands start here...
  offsetpointing 0.1 -10 10 detxy rel
  red go
  offsetpointing -0.3 20 0 detxy rel
  red go
  offsetpointing 0.8 0 -20 detxy rel
  red go
  offsetpointing -1.8 -20 0 detxy rel
  red go

end

warning: guide cam exposure time below lower limit.

spot determination failed after one offset.

Guide star is brighter than limit. R=11.14 --- yes, that's too bright! it was an old script.

Going to run script with a new guide star that is 13.73.
Still we get the warning: guide cam exp time below lower limit, and Geno says that this time the guide camera exptime is even shorter (>~600ms). There are no clouds and the object is not that close to the moon.

Send script with R=15.25 guide star. OK now, but seeing also is a little worse, it seems (1.1-1.4).

Again, note on MODS1r that there is a blotch to the lower right --- around X=170, Y=570-720. Need to check some internal flats from before shutdown and compare with those taken after restart.

05:24 Guide Camera Rotator Center Check

Geno sending on-axis active preset to BS9183 (13th mag--- not complaining).

Stopping guiding

dIE 5"

selectAGW left MODS_L.cfg

readGuideCam left -e 1000

rotator at 0

Slewing rotator to 360 while readGuideCam left -e 90000

/home/lbto/SciPrep/MODS/GuideRotCen/left_guider000003.fits --- this is centered on WFS camera (not exactly).

Trying again ---

Stop guiding, dIE = 5, gprobe 0 0, slew to 360 while taking an exposure.

Sending active preset again.

Stop guiding. readGuideCam left -e 3000 to get image of star in beamsplitter (6, looks like star is in beamsplitter if not perfectly centered in X).

gprobe 0 0

readGuideCam left -e 3000 to get image of star at guidecam_center (looks like 276, 203)

dIE 5"

rotate to 360 and then from 360 to 0. (center looks like 270, 204).

One more time ---

Active preset, stop guiding, readGuideCam left -e 3000 (9). gprobe 0 0, readGuideCam -e 3000 (10).

Geno slewing rotator to 270. no dCA given because we have some other stars in field. (11). slew to -90

dCA = 5". Slewing from -90 to 270. (12... star does not seem to have moved much from center and is rotating on itself).

dCA = -15". Slewing from 270 to -90 (13). Rotator center appears to be at 269 203

dCA=+15, dIE = +5". Slewing from -90 to 270. (14).

dIE = +10" from where we were. Slewing from 270 to -90. Puzzling... the last few rotated images are hugely saturated on BS star, whereas the first ones were not.

06:30 switch to luci.

Field Data @ LFBG

06:58 LUCI1 configured. Geno Set pointing. Send active preset to Stone_P1

07:02 field_collect, Stone_P1.list

07:15 Successfully tested the acquisition of a LUCI1 image from my IDL routines

07:50 Done with field data. 8 points collected in ~0.6" seeing. Will analyze tomorrow

Field Data Analysis

The binodal astigmatism measured is within the measurement error of the analysis compared to the binodal astigmatism measured before shutdown. The measured error of ~ 6" in Tip and ~ 17" in Tilt are small. A change of -10" in Tilt could improve the M1/M2 alignment a bit, but I would say not worth making the change.

 => Field Data Results
             Measured bi-nodal astigmatism
  
                  lfbg
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   4.7", Delta IE =  -4.7"
         Tilt =  15.5", Delta CA = -15.5"
  
  
                  lfbg
              From 0 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   9.4", Delta IE =  -9.4"
         Tilt =  18.7", Delta CA = -18.7"

  • Binodal results from 0 Astigmatism:
    0_astig.jpg

  • Binodal results from 45 Astigmatism:
    45_astig.jpg

LUCI1 Spectra

Longslit Acquisitions and Deep Spectra (not sure if there is a usable MOS mask)

Tell1 for J0100
  • 0019 acq, seeing on guider ~0.5"
  • 0020 slit dX=+1.08"
  • 0021 slit+src
  • 0022 0025 1x60.0s O2DCR 210K spectra (11k ADU peak)
J0100
  • 0026 sky guider FWHM 0.4-0.5"
  • 0027 src
  • 0028 slit dX=+1.65"
  • 0029 src+slit
  • 0030 0037 1x450.0s MER10 210K spectra (stopped to check alignment, src flux is less)
  • 0038 sky, 0039 src back at start. Offset -0.4" in X 0040 after offset, looks good
  • 0041 1x450.0s MER10 210K spectra (AGw warning? Probe can't move to position?)
  • 0042 0043 0044 0045 redo acq.
  • 0046 0059 1x450.0s MER10 210K spectra (seeing went to >1" by end)
  • Seeing got worse in the last half hour
In the first set the source seemed to drift off the slit during/after meridian passage. Is this indicating differential flexure

(or flop) between the AGw and LUCI?

Tell1 for J0100
  • 0060 src
  • 0061 slit (applied +1.14,-1.04" offset while reading out...looks good!)
  • 0062 0065 3x10.0s O2DCR spectra, seeing ~0.45" (counts kind of high)
  • 0066 0069 5x4.0s
K Spectro Flats
  • Telescope@Zenith, lights OFF in dome, CU in
  • 0070 0074 1x7.0s Halo2 lamp OFF
  • 0075 0079 1x7.0s Halo2 lamp ON ? lamp time of 15s in script too short!
  • 0080 0084 1x7.0s Halo2 OFF
  • 0085 0089 1x7.0s Halo2 ON (only ~1800 ADU above background)
  • 0090 0094 1x15.0s Halo2 OFF
  • 0095 0099 1x15.0s Halo2 ON (~10000 ADU per exposure?)

Done @ UT 12:30 (DThompson)

-- %USERSIG{OlgaKuhn - 2014-09-12}%

Comments

 
Topic attachments
I Attachment Action Size Date Who Comment
0_astig.jpgjpg 0_astig.jpg manage 53 K 12 Sep 2014 - 18:23 DougMiller Binodal results from 0 Astigmatism
45_astig.jpgjpg 45_astig.jpg manage 56 K 12 Sep 2014 - 18:48 DougMiller Binodal results from 45 Astigmatism
Topic revision: r8 - 16 Sep 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback